An engineer has designed a valve that will regulate water pressure on an automobile engine. The valve was tested on 180 engines and the mean pressure was 6.2 pounds/square inch (psi). Assume the population standard deviation is 0.8. The engineer designed the valve such that it would produce a mean pressure of 6.3 psi. It is believed that the valve does not perform to the specifications. A level of significance of 0.01 will be used. Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

Solution :

\mu= 6.3

\bar x = 6.2

\sigma = 0.8

n = 180

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (6.2 - 6.3 ) / 0.8 / \sqrt 180

Test statistic = z = -1.68